# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ Abstract
Data assimilation (DA) combines partial observations with a dynamical model to improve state estimation. Filter-based DA uses only past and present data and is the prerequisite for real-time forecasts. Smoother-based DA exploits both past and future observations. It aims to fill in missing data, provide more accurate estimations, and develop high-quality datasets. However, the standard smoothing procedure requires using all historical state estimations, which is storage-demanding, especially for high-dimensional systems. This paper develops an adaptive-lag online smoother for a large class of complex dynamical systems with strong nonlinear and non-Gaussian features, which has important applications to many real-world problems. The adaptive lag allows the DA to utilize only observations within a nearby window, significantly reducing computational storage. Online lag adjustment is essential for tackling turbulent systems, where temporal autocorrelation varies significantly over time due to intermittency, extreme events, and nonlinearity. Based on the uncertainty reduction in the estimated state, an information criterion is developed to systematically determine the adaptive lag. Notably, the mathematical structure of these systems facilitates the use of closed analytic formulae to calculate the online smoother and the adaptive lag, avoiding empirical tunings as in ensemble-based DA methods. The adaptive online smoother is applied to studying three important scientific problems. First, it helps detect online causal relationships between state variables. Second, its advantage of computational storage is illustrated via Lagrangian DA, a high-dimensional nonlinear problem. Finally, the adaptive smoother advances online parameter estimation with partial observations, emphasizing the role of the observed extreme events in accelerating convergence.
@@ 1 Introduction
Complex turbulent nonlinear dynamical systems (CTNDSs) have broad applications across various fields [1, 2, 3, 4]. These systems are characterized by their high dimensionality and multiscale structures, with strong nonlinear interactions occurring between state variables at different spatiotemporal scales. Extreme events, intermittency, and non-Gaussian probability density functions (PDFs) are commonly observed in these systems [5, 6, 7]. State estimation is essential for parameter estimation, prediction, optimal control, and generating complete datasets [8, 9, 10]. However, the turbulent nature of dynamics can amplify small errors in model structure, spatiotemporal solutions, or initial conditions when relying solely on forecasts. Data assimilation (DA), which integrates observations with system dynamics, is widely used to improve state estimation [11, 12, 13, 14, 15]. Given the inevitable uncertainty in state estimation, especially for the unobserved variables in a CTNDS, probabilistic state approaches via Bayesian inference are natural choices. The model provides a prior distribution, while observations inform the likelihood. They are combined to form the posterior distribution for state estimation.
@@
DA can be classified into two categories based on when observational data is incorporated. Filtering uses observations only up to the current time. Serving as the initialization, filter-based state estimation is the prerequisite for real-time forecasts. In contrast, smoothing [15, 16, 17, 18, 19] leverages data from the entire observation period, including future data, which makes it highly effective for optimal state estimation in offline data postprocessing. This helps to fill missing values, minimize bias, and create complete datasets [20]. With the extra information from future observations, smoothing often produces more accurate and less uncertain state estimates than filtering. When the system dynamics and observational mappings are linear, with additive Gaussian noise, the corresponding filtering and smoothing methods are the Kalman filter and the Rauch-Tung-Striebel (RTS) smoother, respectively [16, 21, 22], where the posterior distribution can be computed using closed-form analytical solutions.
@@
Due to the intrinsic nonlinear dynamics and non-Gaussian statistics of CTNDSs, analytic solutions for DA are rarely available. As a result, various numerical and approximate methods have been developed, including the ensemble Kalman filter/smoother and the particle (or sequential Monte Carlo) filter/smoother [14, 15, 23, 24, 25, 26, 27]. These methods are widely used but often face tremendous computational costs, especially in high-dimensional systems [28], which limits the number of particles or ensemble members, potentially causing biases and numerical instabilities [29, 30, 31]. Empirical tuning techniques, such as noise inflation, localization, and resampling, are widely used in practice to mitigate these issues [23, 32, 33, 34]. However, these ad hoc tuning methods are usually quite challenging to implement systematically. Closed-form analytic solutions for DA are thus highly desirable, as they improve computational efficiency, stability, and accuracy, especially in capturing non-Gaussian features, including intermittency and extreme events. They also facilitate theoretical analysis of error and uncertainty in state estimation.
@@
Instead of refining DA schemes directly, computational challenges in state estimation can be addressed by developing approximate models that yield analytic solutions for the posterior distribution. While linear approximations allow for standard methods like the Kalman filter or RTS smoother, linearizing a strongly nonlinear system often leads to biases and instabilities. An alternative is a recently developed class of nonlinear systems that includes many turbulent models in geophysics, fluids, engineering, and neuroscience [35, 36, 37, 38, 39]. Despite their nonlinear dynamics and non-Gaussian statistics, the conditional distributions of unobserved state variables given observations are Gaussian (which is precisely the posterior distribution in the DA context), leading to the term conditional Gaussian nonlinear systems (CGNSs). The CGNS framework allows the use of closed analytic formulae for solving these conditional distributions, helping develop efficient algorithms for filtering, smoothing, and sampling without the ad hoc tuning often needed in ensemble-based DA methods. It also facilitates rigorous analysis of these methods. Additionally, CGNSs have been utilized as surrogate models in various applications, including DA, prediction, preconditioning, and machine learning [39, 40, 41, 42].
@@
The standard smoother-based state estimation procedure involves executing a forward pass for filtering across the entire observational period, followed by a backward pass for smoothing [16, 17]. However, the standard offline smoother requires storing the filter solution for the entire duration before initiating the backward pass, which requires substantial computational storage, particularly in high-dimensional systems. Due to the wide application of smoother-based state estimation, it is of practical importance to develop a computationally efficient and accurate algorithm that significantly reduces storage.
@@
This paper presents a forward-in-time online smoother algorithm with adaptive lags for the CGNS framework, eliminating the need for a full backward pass. The online smoother sequentially updates the current state as new observations become available. By doing so, it effectively addresses the computational storage issue. While online schemes exist for the RTS smoother and ensemble-based methods, the CGNS online smoother has several unique advantages. First, despite the intrinsic nonlinearity of the underlying dynamics, closed analytic formulae are available to compute the nonlinear online smoother. These analytic formulae provide precise and accurate solutions, which avoid numerical and sampling errors as in ensemble-based methods. Second, due to the turbulent nature of the system, observations influence the estimated state only within a short time window, which enhances computational efficiency and reduces storage needs. Third, different from fixed-lag smoothers [17, 27, 43, 44], the lag in the CGNS smoother is adaptively determined. Online lag adjustment is essential for studying turbulent systems, where temporal autocorrelation varies significantly over time due to intermittency, extreme events, and nonlinearity. A fixed lag usually either overuses storage (if the lag is overestimated) or introduces a large bias (if the lag is underestimated). In contrast, an adaptive lag optimizes the use of data and computational storage. Finally, the adaptive lag is systematically determined using an information criterion based on the uncertainty reduction in the posterior distribution [38, 45]. It emphasizes the importance of the posterior uncertainty and differs from some of the existing adaptive lag selection criteria that rely solely on the posterior mean [46]. As closed analytic formulae are available for posterior distributions, the information gain can be computed efficiently and accurately.
@@
The adaptive online smoother for the CGNS framework is applied to studying three important scientific problems. First, the online update of the smoother estimate allows for quantification of the improvement in state estimation by incorporating future information. It facilitates revealing causal dependence between state variables. A nonlinear dyad model with strong non-Gaussian features is utilized for such a study. Second, the CGNS framework is applied to Lagrangian data assimilation (DA), which is a high-dimensional nonlinear problem that has a significant storage requirement [42, 47, 48, 49]. The online smoother allows for the estimation of the unobserved flow states based on Lagrangian tracers. The study highlights the role of the online smoother in reducing computational storage. Finally, the online smoother facilitates developing an online parameter estimation algorithm with partial observations. It helps reveal the role of the observed intermittent extreme events in advancing parameter estimation.
@@
The remainder of this paper is organized as follows. Section 2 introduces the CGNS modeling framework, including the equations for the offline optimal nonlinear filter and smoother state estimation. Section 3 presents the adaptive online smoother. In Section 4, the application of the adaptive online smoother to the three key problems is demonstrated. Section 5 includes the conclusion. The appendices contain detailed analysis and proofs.
2 The Conditional Gaussian Nonlinear System Modelling Framework
@@
Throughout this paper, boldface letters are exclusively used to denote vectors for the sake of mathematical clarity. In this regard, we use lowercase boldface letters to denote column vectors, while Uppercase boldface letters denote matrices. The only exception to this rule is $\mathbf{W}$ (with some subscript or superscript), which denotes a Wiener process. Although in this work, this always corresponds to a column vector, we instead use an uppercase letter due to literary tradition.
@@
Let $t$ denote the time variable, with $t\in[0,T]$, where $T>0$ is allowed to be infinite. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space and $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ to be a filtration of sub-$\sigma$-algebras of $(\Omega,\mathcal{F})$, which we assume is augmented (i.e., complete and right-continuous). For every filtration there exists a smallest such augmented filtration refining $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ (known as its completion), so this is without loss of generality. We let $(\mathbf{x}(t,\omega),\mathbf{y}(t,\omega))$, for $t\in[0,T]$ and $\omega\in\Omega$ (for the rest of this work we drop the event or sample space dependence for notational simplicity, but it is always implied), be a partially observable $(S,\mathcal{A})$-valued stochastic process, where $\mathbf{x}$ is the observable component, while $\mathbf{y}$ is the unobservable component. The theory that follows can be applied mutatis mutandis to any partially observable stochastic process that takes values over a measurable space $(S,\mathcal{A})$ where $S$ is a separable Hilbert space (finite-dimensional or not) over a complete scalar ground field and $\mathcal{A}$ is a $\sigma$-algebra of $S$, but for this work it suffices to consider complex-valued finite-dimensional processes with respect to the usual Euclidean inner product. As such, we let $S=\mathbb{C}^{k+l}$ and $\mathcal{A}=\mathcal{B}_{\mathbb{C}^{k+l}}\equiv\mathcal{B}_{\mathbb{R}^{2(k+l)}}$, with $\mathbf{x}$ being a $k$-dimensional vector and $\mathbf{y}$ an $l$-dimensional one, where $\mathcal{B}_{\mathbb{R}^{2(k+l)}}$ is the Borel $\sigma$-algebra of $\mathbb{R}^{2(k+l)}$, since $\mathrm{dim}_{\mathbb{R}}(\mathbb{C}^{k+l})=2(k+l)$. We assume that
@@
meaning the partially observable random process is (jointly) adapted to the filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$, i.e., for all times $t\in[0,T]$ the random vector defined by $(\mathbf{x}(t,\boldsymbol{\cdot})^{\mathtt{T}},\mathbf{y}(t,\boldsymbol{\cdot})^{\mathtt{T}})^{\mathtt{T}}:\Omega\to S$ is an $(\mathcal{F}_{t};\mathcal{A})$-measurable function. Specifically, this implies that the natural filtration of $\mathcal{F}$ with respect to $\{\mathbf{x}(s)\}_{s\leq t}$, which is the sub-$\sigma$-algebra generated by the observable processes for times $s\leq t$ and is defined by
$\mathcal{F}_{t}^{\mathbf{x}}:=\sigma(\{\mathbf{x}(s)\}_{s\leq t})=\{\mathbf{x}(s)^{-1}[A]=\mathbf{x}(s,\boldsymbol{\cdot})^{-1}[A]:A\in\mathcal{A},\ s\leq t\},$
@@
It is known that, at each time instant $t\in[0,T]$, the optimal estimate in the mean-square sense of some $\mathcal{F}_{t}$-measurable function $h(\mathbf{x}(t),\mathbf{y}(t))$ of $\mathbf{x}$ and $\mathbf{y}$, on the basis of the observations up to time $t$, $\{\mathbf{x}(s)\}_{s\leq t}$, is exactly its conditional expectation conditioned on the observable $\sigma$-algebra at time $t$, $\mathbb{E}\left[\mathbf{h}(\mathbf{x}(t),\mathbf{y}(t))|\mathcal{F}_{t}^{\mathbf{x}}\right]$. This is known as the a-posteriori mean and this assertion of optimality rests on the tacit assumption that $\mathbb{E}\left[\left\|\mathbf{h}(\mathbf{x}(t),\mathbf{y}(t))\right\|_{2}^{2}\right]$ is finite, where $\left\|\boldsymbol{\cdot}\right\|_{2}$ denotes the usual Euclidean norm over $\mathbb{C}^{\mathrm{dim}(\mathbf{h})}$ [35, 36]. Usually $\mathbf{h}$ is a function of the unobserved process $\mathbf{y}$, and in this work we exclusively use $\mathbf{h}=\mathbf{y}$ as to recover the optimal filter and smoother conditional Gaussian statistics of the hidden process when conditioning on the observations.
The goal of optimal state estimation is to characterize the posterior states using a system of stochastic differential equations (SDEs), known as the optimal nonlinear filter or smoother equations, depending on whether we condition on current observations or the entire observational period, respectively. In general, without making specific assumptions about the structure of the processes $\mathbf{h}$ and $\mathbf{x}$, determining $\mathbb{E}\left[\mathbf{h}|\mathcal{F}_{t}^{\mathbf{x}}\right]$ is challenging. Nevertheless, the conditional Gaussian nonlinear system (CGNS) has a key advantage in its conditional Gaussianity of the posterior distributions, which can be written down using closed analytic formulae. Note that the classical Kalman-Bucy filter [21] is the simplest example of a CGNS [36]. Despite the conditional Gaussianity, these coupled systems remain highly nonlinear and the associated marginal and joint distributions are highly non-Gaussian, which allows the systems to capture many realistic turbulent features.

### v2

@@ Abstract
Data assimilation (DA) combines partial observations with dynamical models to improve state estimation. Filter-based DA uses only past and present data and is the prerequisite for real-time forecasts. Smoother-based DA exploits both past and future observations. It aims to fill in missing data, provide more accurate estimations, and develop high-quality datasets. However, the standard smoothing procedure requires using all historical state estimations, which is storage-demanding, especially for high-dimensional systems. This paper develops an adaptive-lag online smoother for a large class of complex dynamical systems with strong nonlinear and non-Gaussian features, which has important applications to many real-world problems. The adaptive lag allows the utilization of observations only within a nearby window, thus reducing computational complexity and storage needs. Online lag adjustment is essential for tackling turbulent systems, where temporal autocorrelation varies significantly over time due to intermittency, extreme events, and nonlinearity. Based on the uncertainty reduction in the estimated state, an information criterion is developed to systematically determine the adaptive lag. Notably, the mathematical structure of these systems facilitates the use of closed analytic formulae to calculate the online smoother and adaptive lag, avoiding empirical tunings as in ensemble-based DA methods. The adaptive online smoother is applied to studying three important scientific problems. First, it helps detect online causal relationships between state variables. Second, the advantage of reduced computational storage expenditure is illustrated via Lagrangian DA, a high-dimensional nonlinear problem. Finally, the adaptive smoother advances online parameter estimation with partial observations, emphasizing the role of the observed extreme events in accelerating convergence.
@@ 1 Introduction
Complex turbulent nonlinear dynamical systems (CTNDSs) have broad applications across various fields [1, 2, 3]. These systems are characterized by their high dimensionality and multiscale structures, with strong nonlinear interactions occurring between state variables at different spatiotemporal scales. Extreme events, intermittency, and non-Gaussian probability density functions (PDFs) are also commonly observed [4, 5, 6].
State estimation in CTNDSs is essential for parameter estimation, prediction, optimal control, and generating complete datasets [7, 8, 9]. However, the turbulent nature of the dynamics can amplify small errors in the model structure, spatiotemporal solutions, or initial conditions when relying solely on forecasts. Data assimilation (DA), which integrates observations with system dynamics, is widely used to improve state estimation [10, 11, 12, 13, 14]. Given the inevitable uncertainty in state estimation, especially for the unobserved variables of CTNDSs, probabilistic approaches via Bayesian inference are natural choices. The model provides a prior distribution, while observations inform the likelihood. These are then combined to form the posterior distribution for state estimation.
@@
DA can be classified into two categories, based on when the observational data are incorporated. Filtering uses observations only up to the current time [12, 13, 14]. Serving as the initialization, filter-based state estimation is a prerequisite for real-time forecasts. In contrast, smoothing leverages data from the entire observation period [14, 15, 16, 17], including future data, which makes it highly effective for optimal state estimation in offline data postprocessing. This helps to fill in missing values, minimize bias, and create complete datasets [18]. With the extra information from future observations, smoothing often produces more accurate and less uncertain state estimates than filtering. When the system dynamics and observational mappings are linear, with additive Gaussian noise, the corresponding filtering and smoothing methods are the Kalman(–Bucy) filter and the Rauch–Tung–Striebel (RTS) smoother, respectively [15, 19, 20], where the posterior distribution is Gaussian and can be computed using closed-form analytical solutions.
@@
Due to the intrinsic nonlinear dynamics and non-Gaussian statistics of CTNDSs, analytic solutions for DA are rarely available. As a result, various numerical and approximate methods have been developed, including the ensemble Kalman filter/smoother and the particle (or sequential Monte Carlo) filter/smoother [13, 14, 21, 22, 23, 24]. These methods are widely used but often face tremendous computational costs, especially in high-dimensional systems [25], which limits the number of particles or ensemble members, potentially causing biases and numerical instabilities [26, 27, 28]. Empirical tuning techniques, such as noise inflation, localization, and resampling, are widely used in practice to mitigate these issues [21, 29, 30, 31]. However, these ad hoc tuning methods are usually quite challenging to implement systematically. Closed-form analytic solutions for DA are thus highly desirable, as they improve computational efficiency, stability, and accuracy, especially in capturing non-Gaussian features, including intermittency and extreme events. They also facilitate the theoretical analysis of the error and uncertainty during state estimation.
@@
Instead of refining DA schemes directly, computational challenges in state estimation can be addressed by developing approximate models that yield analytic solutions for the posterior distribution. While linear approximations allow for standard methods like the Kalman filter or RTS smoother, linearizing a strongly nonlinear system often leads to biases and instabilities. An alternative is a recently developed class of nonlinear systems that includes many turbulent models in geophysics, fluids, engineering, and neuroscience [32, 33, 34]. Despite their nonlinear dynamics and non-Gaussian statistics, the conditional distributions of the unobserved state variables given the observations, which are precisely the posterior distributions in the DA context, are Gaussian, leading to the term conditional Gaussian nonlinear systems (CGNSs). The CGNS framework allows the use of closed analytic formulae for solving these conditional distributions, helping develop efficient algorithms for filtering, smoothing, and sampling without the ad hoc tuning often needed in ensemble-based DA methods. CGNSs have been utilized as surrogate models in various applications, including DA, prediction, preconditioning, and machine learning [35, 36, 37].
@@
The standard smoother-based state estimation procedure involves executing a forward pass for filtering across the entire observational period, followed by a backward pass for smoothing [15, 16]. However, the standard offline smoother requires storing the filter solution for the entire duration before initiating the backward pass, which requires substantial computational storage, particularly in high-dimensional systems. Due to the wide application of smoother-based state estimation, it is of practical importance to develop a computationally efficient and accurate algorithm that has the potential to significantly reduce storage requirements.
@@
This paper develops a forward-in-time online smoother algorithm with adaptive lags for the CGNS framework, eliminating the need for a full backward pass. The online smoother sequentially updates the current state as new observations become available. By doing so, it effectively addresses the computational storage issue. While online schemes exist for the RTS smoother and ensemble-based methods, the CGNS online smoother has several unique advantages. First, despite the intrinsic nonlinearity of the underlying dynamics, closed analytic formulae are available to compute the nonlinear online smoother, providing precise and accurate solutions which avoid numerical and sampling errors as in ensemble-based methods. Second, due to the turbulent nature of the system, observations influence the estimated state only within a short time window, which enhances computational efficiency and reduces storage needs. Third, different from fixed-lag smoothers [24, 38, 39], the lag in the CGNS smoother is adaptively determined. Online lag adjustment is essential for studying turbulent systems, where temporal autocorrelation varies significantly over time due to intermittency, extreme events, and nonlinearity. A fixed lag usually either overuses storage (when overestimated) or introduces a large bias (if underestimated). In contrast, an adaptive lag implicitly optimizes the use of data and computational storage. Finally, the adaptive lag is systematically determined using an information criterion based on the uncertainty reduction in the posterior distribution. It emphasizes the importance of the posterior uncertainty and differs from some of the existing adaptive lag selection criteria that rely solely on the posterior mean [40]. As closed analytic formulae are available for posterior distributions, the information gain can be computed efficiently and accurately.
@@
The adaptive online smoother for CGNSs is employed to study three important scientific problems. First, the online update of the smoother estimate allows for quantification of the improvement in state estimation by incorporating future information. This facilitates the inference of causal relationships between the state variables. A nonlinear dyad model with strong non-Gaussian features is utilized for such a study. Second, the CGNS framework is applied to Lagrangian data assimilation, which is a high-dimensional nonlinear problem that has a significant storage requirement [41, 42, 43]. The online smoother allows for the estimation of the unobserved flow states based on Lagrangian tracers. This study highlights the role of the adaptive-lag online smoother in reducing computational storage needs, when compared to its fixed-lag variant. Finally, the online smoother facilitates developing an online parameter estimation algorithm with partial observations. It helps reveal the role of the observed intermittent extreme events in advancing parameter estimation.
@@
The remainder of this paper is organized as follows. Section 2 introduces the CGNS modeling framework, including the equations for the optimal nonlinear filter and offline smoother state estimation. Section 3 presents the adaptive online smoother. In Section 4, the application of the adaptive online smoother to the three key problems is demonstrated. Section 5 includes the conclusion. The appendices contain the analysis and proofs, as well as additional details.
2 The Conditional Gaussian Nonlinear System Modeling Framework
@@
Throughout this paper, boldface letters are exclusively used to denote vectors for the sake of mathematical clarity. In this regard, we use lowercase boldface letters to denote column vectors, while Uppercase boldface letters denote matrices. The only exception to this rule is $\mathbf{W}$ (with some subscript or superscript), which denotes a Wiener process. Although in this work this always corresponds to a column vector, we instead use an uppercase letter due to literary tradition.
@@
Let $t$ denote the time variable, with $t\in[0,T]$, where $T>0$ is allowed to be infinite. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space and $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ to be a filtration of sub-$\sigma$-algebras of $(\Omega,\mathcal{F})$, which we assume is augmented (i.e., complete and right-continuous). For every filtration there exists a smallest such augmented filtration refining $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ (known as its completion), so this is without loss of generality. We let $\big(\mathbf{x}(t,\omega),\mathbf{y}(t,\omega)\big)$, for $t\in[0,T]$ and $\omega\in\Omega$ (for the rest of this work we drop the event or sample space dependence for notational simplicity, but it is always implied), be a partially observable $(S,\mathcal{A})$-valued stochastic process, where $\mathbf{x}$ is the observable component, while $\mathbf{y}$ is the unobservable component. The theory that follows can be applied mutatis mutandis to any partially observable stochastic process that takes values over a measurable space $(S,\mathcal{A})$ where $S$ is a separable Hilbert space (finite-dimensional or not) over a complete scalar ground field and $\mathcal{A}$ is a $\sigma$-algebra of $S$, but for this work it suffices to consider complex-valued finite-dimensional processes with respect to the usual Euclidean inner product. As such, we let $S=\mathbb{C}^{k+l}$ and $\mathcal{A}=\mathcal{B}_{\mathbb{C}^{k+l}}\equiv\mathcal{B}_{\mathbb{R}^{2(k+l)}}$, with $\mathbf{x}$ being a $k$-dimensional vector and $\mathbf{y}$ an $l$-dimensional one, where $\mathcal{B}_{\mathbb{R}^{2(k+l)}}$ is the Borel $\sigma$-algebra of $\mathbb{R}^{2(k+l)}$, since $\mathrm{dim}_{\mathbb{R}}(\mathbb{C}^{k+l})=2(k+l)$. We assume that
@@
meaning the partially observable random process is (jointly) adapted to the filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$, i.e., for all times $t\in[0,T]$ the random vector defined by $\big(\mathbf{x}(t,\cdot)^{\mathtt{T}},\mathbf{y}(t,\cdot)^{\mathtt{T}}\big)^{\mathtt{T}}:\Omega\to S$ is an $(\mathcal{F}_{t};\mathcal{A})$-measurable function. Specifically, this implies that the natural filtration of $\mathcal{F}$ with respect to $\{\mathbf{x}(s)\}_{s\leq t}$, which is the sub-$\sigma$-algebra generated by the observable processes for times $s\leq t$ and is defined by
$\mathcal{F}_{t}^{\mathbf{x}}:=\sigma\big(\{\mathbf{x}(s)\}_{s\leq t}\big)=\big\{\mathbf{x}(s)^{-1}[A]=\mathbf{x}(s,\cdot)^{-1}[A]:A\in\mathcal{A},\ s\leq t\big\},$
@@
It is known that, at each time instant $t\in[0,T]$, the optimal estimate in the mean-square sense of some measurable function $\mathbf{h}(t,\mathbf{x},\mathbf{y})$ on the basis of the observations up to time $T^{\prime}\in[t,T]$, $\{\mathbf{x}(s)\}_{s\leq T^{\prime}}$, is exactly its conditional expectation conditioned on the observable $\sigma$-algebra at time $T^{\prime}$, $\mathbb{E}\big[\mathbf{h}(t,\mathbf{x},\mathbf{y})|\mathcal{F}_{T^{\prime}}^{\mathbf{x}}\big]$. This is known as the a-posteriori mean and this assertion of optimality rests on the tacit assumption that $\mathbb{E}\big[\left\|\mathbf{h}(\mathbf{x},\mathbf{y})\right\|_{2}^{2}\big]$ is finite, where $\left\|\cdot\right\|_{2}$ denotes the usual Euclidean norm over $\mathbb{C}^{\mathrm{dim}(\mathbf{h})}$ [32, 33]. Usually $\mathbf{h}$ is a function of the unobserved process $\mathbf{y}$ and in this work we exclusively use $\mathbf{h}=\mathbf{y}$ as to recover the optimal filter and smoother conditional statistics of the hidden process when conditioning on the observations.
The goal of optimal state estimation is to characterize the posterior states using a system of stochastic differential equations (SDEs), known as the optimal nonlinear filter or smoother equations depending on whether we condition on current observations or the entire observational period, respectively. In general, without making specific assumptions about the structure of the processes $\mathbf{h}$ and $\mathbf{x}$, determining $\mathbb{E}\left[\mathbf{h}|\mathcal{F}_{T^{\prime}}^{\mathbf{x}}\right]$ is challenging. CGNSs resolve this by having the key advantage of conditional Gaussian posterior distributions, which can be written down using closed analytic formulae. The linear state-observation system filtered by the classical Kalman–Bucy filter [19, 20] is the simplest example of a CGNS [33, 44]. Despite the conditional Gaussianity, these coupled systems remain highly nonlinear, with the associated marginal and joint distributions being highly non-Gaussian, which allows the systems to capture many realistic turbulent features.